OpenStudy (fibonaccichick666):

Just need this one checked. Solve \[y'''-2y'+y=te^t\] Char eq. \[r^3-2r+1=0\]\[(r-1)(r^2+r-1)=0\]\[r=1,~-.5(1+\sqrt 5),~ .5(\sqrt 5-1)\]

OpenStudy (fibonaccichick666):

then my homogeneous solution is\[ y= c_1e^t+c_2e^{-.5t(1+\sqrt 5)}+c_3e^{.5t(\sqrt5-1)}\]

OpenStudy (fibonaccichick666):

Do you agree to this point?

OpenStudy (dan815):

use undetermined coeff of variation of paramets for the particular solution

OpenStudy (dan815):

or* variation of parameters

OpenStudy (fibonaccichick666):

but do you agree to that point?

OpenStudy (fibonaccichick666):

I din't muff anything up through there?

OpenStudy (fibonaccichick666):

ok, so for MOUC \(y_p(t)=(t-1)e^t\)

OpenStudy (fibonaccichick666):

yea?

OpenStudy (fibonaccichick666):

@amistre64 do you mind checking this one so far?

OpenStudy (fibonaccichick666):

Apparently I am wrong

OpenStudy (fibonaccichick666):

the partiticular sln should be of the form (At^2+Bt)e^t

OpenStudy (amistre64):

\[y'''-2y'+y=te^t\] \[(e^{rx})'''-2(e^{rx})'+(e^{rx})=0\] \[r^3(e^{rx})-2r(e^{rx})+(e^{rx})=0\] \[r^3-2r+1=0\] \[r=1,\frac12(-1\pm\sqrt{5}) \] \[y_h=c_0exp(x)+c_1exp(\frac{-1-\sqrt{5}}{2}x)+c_2exp(\frac{-1+\sqrt{5}}{2}x)\]

OpenStudy (amistre64):

x .. t... meh

OpenStudy (fibonaccichick666):

it's ok,

OpenStudy (fibonaccichick666):

I got what you meant

OpenStudy (amistre64):

ever work a wronskian?

OpenStudy (fibonaccichick666):

yea, but I don't have two solutions

OpenStudy (fibonaccichick666):

ohhhhhhhH!H!!H!HH!H!!HH!H!!

OpenStudy (amistre64):

you have a degree 3 and 3 soluitions

OpenStudy (fibonaccichick666):

ce^t and ce^{.5(sqrt5-1)t

OpenStudy (fibonaccichick666):

that just skipped over my head. That will be faster. Variation of parameters it is

OpenStudy (fibonaccichick666):

thanks!

OpenStudy (amistre64):

itll be a 3x4 matrix i believe. good luck :)

OpenStudy (fibonaccichick666):

nah, we only need two of em

OpenStudy (fibonaccichick666):

right?

OpenStudy (amistre64):

my memory says no. if we has a 2nd order y'' then 2 yes we have a y''' order so its a 3x3 setup

OpenStudy (amistre64):

we have 3 linearly independant terms in the homogenous

OpenStudy (fibonaccichick666):

ew. Then I feel MOUC would be better in this instance

OpenStudy (amistre64):

you dont know the 3x3 determinant shortcut? pretty much reduces to the 2x2 shortcut in a way; multiplicaiton of diags

OpenStudy (fibonaccichick666):

I know, but it's not pretty and a mess to integrate

OpenStudy (amistre64):

ok, ive just never come across MOUS before so i wouldnt know how that operates

OpenStudy (fibonaccichick666):

method of undertermined coefficients

OpenStudy (amistre64):

i call that variation of parameters ... of course my vocab in this area migh tbe rusty

OpenStudy (fibonaccichick666):

variation of parameters uses integrals and wronskian

OpenStudy (fibonaccichick666):

the answer to this is ridiculous

OpenStudy (amistre64):

\[y_p=Af+Bg+Ch\] \[y'_p=Af'+Bg'+Ch'+(A'f+B'g+C'h:=0)\] \[y''_p=Af''+Bg''+Ch''+(A'f'+B'g'+C'h':=0)\] \[y'''_p=Af'''+Bg'''+Ch'''+(A'f''+B'g''+C'h'':=te^t)\] waaaaiiiittt a minute, we have constant coefficients so isnt this a bit much for it? even tho this should still pan out ...

OpenStudy (fibonaccichick666):

With initial conditions

OpenStudy (amistre64):

i just pooed myself :) Laplace transform maybe? since we have all those ynots

OpenStudy (fibonaccichick666):

that may work better

OpenStudy (fibonaccichick666):

that will work better :D

OpenStudy (amistre64):

you didnt include the IVP stuff to start with or i might have suggested that ... well about the same time but still...

OpenStudy (fibonaccichick666):

well, I shall do it now

OpenStudy (fibonaccichick666):

so \[L(y)=\frac{1}{(s-1)^2(s^3-2s+1)}\]

OpenStudy (fibonaccichick666):

yea?

OpenStudy (fibonaccichick666):

so then \[L(t)=\frac{1}{(s-1)^3} *\frac{1}{s^2+s-1} \]

OpenStudy (amistre64):

\[s^3 L\{y\}+L\{y\}-2sL\{y\}=\frac{1}{(s-1)^2}\]

OpenStudy (amistre64):

you seem to be doing fine :)

OpenStudy (fibonaccichick666):

My issue is how to do the inverse L(t)

OpenStudy (amistre64):

complete the squuare on the s^2+s+1 .... and i think your going to decomp into a sum of fraction right?

OpenStudy (fibonaccichick666):

uhm, I know the roots of that, but the issue is how do I do the inverse L(t) for 3 things multiplied together

OpenStudy (fibonaccichick666):

like this is the final answer http://www.wolframalpha.com/input/?i=inverse+Laplace+transform+%281%2F%5B%28s-1%29%5E3%28s%5E2%2Bs-1%29%5D%29

OpenStudy (amistre64):

thats definiantly a better siplification than the other output

OpenStudy (fibonaccichick666):

yea, but how do I get it....

OpenStudy (amistre64):

\[\frac{1}{(s-1)^3(s^2+s+1)}=\frac{As^2+Bs+C}{(s-1)^3}+\frac{Ms+N}{(s^2+s+1)}\]

OpenStudy (fibonaccichick666):

ah

OpenStudy (fibonaccichick666):

thank you

OpenStudy (amistre64):

yw

OpenStudy (fibonaccichick666):

I got it from here now